OpenStudy (anonymous):

Can You Tell Me If My Answers Are Correct! 1. 10(x+2)+3(5x+4)=-18 Answer: x= 16 2. -17x9-3x-8+9x=-10 Answer: x=-1 3. 3-4^2*5/ 8+3 Answer: -4 4. 8x-17=-57 Answer: x= -5 5. x/-12+4=-7 Answer: x=132

OpenStudy (cwrw238):

1. incorrect - if you plug 16 in for x the total comes to 432

OpenStudy (cwrw238):

10(x+2)+3(5x+4)=-18 10x + 20 + 15x + 12 = -18 25x = -20-12-18 25x = -50

OpenStudy (cwrw238):

can you understand what i just did?

OpenStudy (anonymous):

Yea! I Just Made A Small Mistake! But Is Everything Else Good?

OpenStudy (cwrw238):

4 and 5 are ok

OpenStudy (anonymous):

2 Is Wrong?

OpenStudy (cwrw238):

not sure what 3 is exactly is it 3 - 4^2 * 5 --------- 9 + 3

OpenStudy (cwrw238):

* i meant 8+3 on the bottom

OpenStudy (anonymous):

3- 4^2*5 Divided By 8+3

OpenStudy (anonymous):

\[3-4^2*5\div8+3\]

OpenStudy (cwrw238):

ok that 3 - 16*5 -77 ------- = -- = -7 11 11

OpenStudy (cwrw238):

-17x9-3x-8+9x=-10 what is the first term -17x9 mean?

OpenStudy (anonymous):

Wouldnt You Add The Like Terms?

OpenStudy (anonymous):

He Has To Be! Negative 4

OpenStudy (anonymous):

I Mean It!

OpenStudy (cwrw238):

i dont understand what -17x9 is is it -17 times 9 or -17 times x times 9 maybe

OpenStudy (cwrw238):

what which question must be negative 4?

OpenStudy (anonymous):

Sorry Typed The Problem Wrong! \[-17x+9-3x-8+9x=-10\]

OpenStudy (anonymous):

Nvm! I Got Confused With The Answer( Negative 4 Thing)

OpenStudy (cwrw238):

oh ok - noproblem lets plug in your answer x = -1 into the equation -17*-1 + 9 -3*-1 - 8 + 9*-1 = 17 + 9 + 3 - 8 - 9 = 26 +3 - 17 = 12 so theres something wrong there

OpenStudy (cwrw238):

−17x+9−3x−8+9x=−10 bring like terms together -17x - 3x + 9x + 9 - 8 = -10 -17x - 3x + 9x + 1 = -10 now subtract 1 from both sides of equation -17x - 3x + 9x = -11 -20x + 9x = -11 -11x = -11 x = 1 x = 1, you were close - just got the sign wrong